OpenStudy (anonymous):

HELP NEEDED: The number of yeast cells in a laboratory culture increases rapidly initially but levels off eventually. The population is modeled by the function n = f(t) = a/(1 + be−0.9^t) where t is measured in hours. At time t = 0 the population is 20 cells and is increasing at a rate of 16 cells/hour. Find the values of a and b.

OpenStudy (anonymous):

you need two equations to solve for two unknowns a and b Solve for f(0) and f(1) then find a and b with this system of equations

OpenStudy (anonymous):

Hey, I cant seem to find the two equations

OpenStudy (joannablackwelder):

When t=0, n=20 and when t=1, n=36.

OpenStudy (joannablackwelder):

Plug in values to make 2 equations and solve the system to find a and b.

OpenStudy (anonymous):

where is t1, n=36 coming from

OpenStudy (joannablackwelder):

Since, the population is increasing at 16 cells per hr.

OpenStudy (joannablackwelder):

We were given that n=20 at t=0, so at t=1, n=20+16.

OpenStudy (joannablackwelder):

We could also say that when t=3, n=20+16+16 but since we only have 2 constants to fine, we only need 2 equations.

OpenStudy (anonymous):

I see the logic but do you get the same answer if you get two equations f(0) = 20 and f'(0) = 16 since rate of change is first derivative

OpenStudy (joannablackwelder):

I think that should work too.

OpenStudy (joannablackwelder):

Is the function n=a/(1+be^[-0.9t])?

OpenStudy (anonymous):

that looks correct

OpenStudy (joannablackwelder):

I got a=79.66 and b=2.98

OpenStudy (joannablackwelder):

Using the system of equations method.

OpenStudy (anonymous):

is it the right answer?

OpenStudy (joannablackwelder):

I am working it again with the derivative method to check it.

OpenStudy (anonymous):

20 = a/(1+b) 16 = 9ab/(1+b)^2

OpenStudy (anonymous):

.9

OpenStudy (joannablackwelder):

Yes, those are the equations I got using the derivative method.

OpenStudy (joannablackwelder):

Solving those equations simultaneously, I got a=180, b=8. Let me check my work. Just a minute.

OpenStudy (anonymous):

Hi guys, how did the second equation ?

OpenStudy (anonymous):

20 = a/(1+b) 16 = .9ab/(1+b)^2

OpenStudy (anonymous):

f'(0)

OpenStudy (anonymous):

can you show me?

OpenStudy (joannablackwelder):

Neither answer that I got is a function that starts at 20, increases rapidly, then levels off. I think I may have misunderstood the original equation.

OpenStudy (joannablackwelder):

To get the second equation, you use the quotient rule and plug in 0 for t.

OpenStudy (joannablackwelder):

@bobobobobb , do you know how to calculate a derivative using the quotient rule?

OpenStudy (anonymous):

Yes, are you just diffrentiating a/1+b?

OpenStudy (anonymous):

did you use quadratic formula to find b

OpenStudy (joannablackwelder):

No, I differentiated the original equation.

OpenStudy (anonymous):

when you take the first derivative plug in t =0 you end up with the second equation if you make a the subject in the first equation and plug it in in the second equation, you end with a quadratic equation to solve for b